# SIDA: Social Media Image Deepfake Detection, Localization and Explanation with Large Multimodal Model

### v1

@@ 5 Experiments
Implementation Details. We choose LISA as the base large vision language model due to its strong capability for reasoning-based localization. We fine-tuned both LISA-7B-v1 and LISA-13B-v1 on the SID-Set using LoRA, setting $\alpha$ at 16 and the dropout rate at 0.05. The input images are resized to $1024\times 1024$. The loss weights in Eq. (6) for the detection ($\lambda_{\text{det}}$), the text generation ($\lambda_{\text{txt}}$), and the localization ($\lambda_{\text{mask}}$) are set to 1.0, respectively. The localization loss weights in Eq. (5) for the $\lambda_{\text{bce}}$ and $\lambda_{\text{dice}}$ are set to 2.0 and 0.5, respectively. To determine the optimal weight configuration, we perform ablation studies as detailed in Section 5.5. During the detection and localization training stage, the image encoder is frozen, and all other modules are trainable. For the text generation stage, only vision-language models are fine-tuned using the LoRA strategy. The initial learning rate is set to $1\times 10^{-4}$, with a batch size of 2 per device and a gradient accumulation step of 10. We use two NVIDIA A100 GPUs (40GB each). Training for SIDA-7B and SIDA-13B took 48 hours and 72 hours, respectively.
Table 2: Comparison of SIDA with other deepfake detection methods. Values outside parentheses are evaluated with open-source models directly on the SID-Set; values in parentheses indicate performance changes after training with the SID-Set. The best results are in bold.
@@ 5.3 Robustness Study
JPEG 70 94.4 95.1 92.2 71.8 44.3
JPEG 80 93.7 94.5 91.8 71.1 43.7
Resize 0.5 94.3 96.1 92.8 72.5 45.2
Resize 0.75 94.9 96.6 93.1 73.0 45.5
Gaussian 10 91.9 94.3 89.1 70.2 43.0
Gaussian 5 93.4 94.9 90.3 71.0 43.5
SIDA-7B 97.3 97.5 94.6 73.9 45.8
@@ Appendix D Detailed DataSet Creation Process
Tampered Images. We detail each step of the tampered image generation process. Two separate directories were set up: one for object replacement, where entire objects (e.g., animals, vehicles, household items) are swapped with similar classes to generate new scenarios, and another for attribute replacement, which modifies specific features or characteristics of objects (e.g., changing an animal’s emotion or activity, such as making a “dog” appear “happy”or “running”). Figures 18 and 19 illustrate the detailed directories for object and attribute replacements, respectively. Additionally, we used GPT-4o to generate ground truth descriptions for three types of inputs: (1) real images accompanied by prompts, (2) fully synthetic images with corresponding prompts, and (3) tampered images along with prompts and their respective tampered masks, as illustrated in Figure 7. The Figure 3 details each step involved in generating tampered images using the replacement strategies.
Figure 8: Visual comparison of SIDA with other localization methods. Both MVSS-Net and LISA were fine-tuned on the SID-Set for this evaluation.

### v2

@@ 5 Experiments
Implementation Details. We choose LISA as the base large vision language model due to its strong capability for reasoning-based localization. We fine-tuned both LISA-7B-v1 and LISA-13B-v1 on the SID-Set using LoRA, setting $\alpha$ at 16 and the dropout rate at 0.05. The input images are resized to $1024\times 1024$. The loss weights in Eq. (6) for the detection ($\lambda_{\text{det}}$), the text generation ($\lambda_{\text{txt}}$), and the localization ($\lambda_{\text{mask}}$) are set to 1.0, respectively. The localization loss weights in Eq. (5) for the $\lambda_{\text{bce}}$ and $\lambda_{\text{dice}}$ are set to 1.0. To determine the optimal weight configuration, we perform ablation studies as detailed in Section 5.5. During the detection and localization training stage, the image encoder is frozen, and all other modules are trainable. For the text generation stage, only vision-language models are fine-tuned using the LoRA strategy. The initial learning rate is set to $1\times 10^{-4}$, with a batch size of 2 per device and a gradient accumulation step of 10. We use two NVIDIA A100 GPUs (40GB each). Training for SIDA-7B and SIDA-13B took 48 hours and 72 hours, respectively. We divided the SID-Set into a training set, validation set, and test set using a 7:1:2 ratio.
Table 2: Comparison of SIDA with other deepfake detection methods. Values outside parentheses are evaluated with open-source models directly on the SID-Set; values in parentheses indicate performance changes after fine-tuning the models using the SID-Set on the validation set. ${{\color[rgb]{1,0,0}Red}}$ indicates improvement, and ${{\color[rgb]{0,1,0}green}}$ indicates a decrease in performance. The best results are in bold.
@@ 5.3 Robustness Study
JPEG 70 89.4 90.1 86.2 71.8 42.3
JPEG 80 88.7 89.5 85.8 71.1 41.7
Resize 0.5 89.3 91.1 86.8 72.5 43.2
Resize 0.75 89.9 91.6 87.1 73.0 43.5
Gaussian 10 86.9 89.3 84.1 70.2 41.0
Gaussian 5 88.4 89.9 85.3 71.0 41.5
SIDA-7B 93.5 93.5 87.3 73.9 43.8
@@ Appendix B Detailed Comparison
Figure 8: Visual comparison of SIDA with other localization methods. Both approaches were fine-tuned on the SID-Set for this evaluation.
@@ Appendix D Detailed DataSet Creation Process
Tampered Images. We detail each step of the tampered image generation process. Two separate directories were set up: one for object replacement, where entire objects (e.g., animals, vehicles, household items) are swapped with similar classes to generate new scenarios, and another for attribute replacement, which modifies specific features or characteristics of objects (e.g., changing an animal’s emotion or activity, such as making a “dog” appear “happy”or “running”). Figures 18 and 19 illustrate the detailed directories for object and attribute replacements, respectively. Furthermore, we employed GPT-4o to generate ground truth descriptions for three distinct input types: (1) real images paired with prompts, (2) fully synthetic images accompanied by corresponding prompts, and (3) tampered images provided with both prompts and their associated tampered masks, as shown in Figure 7. Figure 3 elaborates on each step of the tampered image generation process using these replacement strategies. To further enrich the diversity of SID-Set, we integrated image segments from Magicbrush [82]. Additional information about Magicbrush is available on its project website [^6] https://osu-nlp-group.github.io/MagicBrush/.
Appendix E Experts and Human Evaluation
We engaged five experts to undertake the following three tasks:
Model Selection: The experts examined approximately 1,000 images to assess the consistency and quality of the generated outputs. Based on their evaluations, we selected FLUX and latent-diffusion as our default generative models due to their superior performance.
Image Quality Assessment: Following image generation, the experts evaluated the realism of the outputs. To standardize this process, we introduced a five-point rating scale for image realism, ranging from 0 (lowest quality) to 5 (highest quality). Images scoring below 3 were flagged as unnatural or defective. These flagged images underwent a secondary review by the experts, after which all identified flawed images were excluded to maintain the dataset’s overall quality.
Textual Description Evaluation: The experts systematically reviewed 3,000 textual descriptions produced by GPT-4 to verify their semantic accuracy and alignment with the corresponding images. This evaluation adhered to three key criteria: (1) Accuracy – ensuring the description accurately reflects the image’s visual content; (2) Clarity – confirming the description is concise, unambiguous, and easily comprehensible; and (3) Consistency – verifying coherence with similar prompts or scenarios across the dataset.
These refinements have been incorporated into our revised approach. We appreciate the constructive feedback and valuable guidance provided, which have significantly strengthened this process.
@@ Appendix E Experts and Human Evaluation
Acknowledgments
This work is supported by The Alan Turing Institute (UK) through the project ‘Turing-DSO Labs Singapore Collaboration’ (SDCfP2\100009).